OpenStudy (anonymous):

Explain, in complete sentences and in your own words, the answers to the following questions relating to complex numbers. What are the characteristics of a complex number? What is the relationship between a complex number and its conjugate? Describe the usefulness of the conjugate and its effect on other complex numbers.

OpenStudy (richyw):

if \(z=a+ib\) where \(a,b\in\mathbb{R}\) then \(\bar{z}=a-ib\)

OpenStudy (anonymous):

A complex number must have a real and imaginary part. It can be in the form: a + bi Where a and b are real numbers and i is the standard imaginary unit i^2 = -1. The relationship is that the product of a complex number and it's respective complex conjugate is a real number. Conjugation of a complex number describes an axial symmetry of the complex plane. To conjugate a complex number, reflect its position through the real axis.